An ideal gas in a sealed container has an initial volume of 2.45 L. At constant pressure, it is cooled to 15.00 ∘C, where its final volume is 1.75 L. What was the initial temperature?
1
Expert's answer
2020-10-19T14:28:13-0400

"\frac{V_1}{V_2}=\frac{T_1}{T_2}"


Given that; "V_1=2.45L, V-2=1.75L, T_2=19^\degree C=273+15=288^\degree K"

"T1=?"


"\frac{V_1}{V_2}=\frac{T_1}{T_2}"


"T_1=(\frac{2.45}{1.75})\times288=403.2^\degree K"


"=403.2-273=130.2^\degree C"